OpenStudy (anonymous):

(4x^4-4x^3+5x^2-4x+1)/(2x-1)

OpenStudy (asnaseer):

is this a remainder problem again?

OpenStudy (anonymous):

problley

OpenStudy (asnaseer):

you can use the same theorem here by rearranging the equation as follows:\[\frac{4x^4-4x^3+5x^2-4x+1}{2x-1}=\frac{4x^4-4x^3+5x^2-4x+1}{2(x-0.5)}\]

OpenStudy (asnaseer):

so now your f(x) is given by:\[f(x)=\frac{4x^4-4x^3+5x^2-4x+1}{2}\]and the remainder would be calculated by working out f(0.5)

OpenStudy (anonymous):

so i divide all that by 2

OpenStudy (asnaseer):

substitute x=0.5 into f(x) and the resulting value will be your remainder

OpenStudy (anonymous):

i dnt get it!!!!

OpenStudy (asnaseer):

i.e. work this out:\[\frac{4(0.5)^4-4(0.5)^3+5(0.5)^2-4(0.5)+1}{2}\]here I have replaced all occurrences of x in the expression for f(x) by the value 0.5

OpenStudy (asnaseer):

I would sugest you study this to become more familiar with it: http://www.mathsisfun.com/algebra/polynomials-remainder-factor.html the remainder theorem is described about a quarter of the way down on that web page.

OpenStudy (anonymous):

i still dnt get it

sam (.sam.):

Or just use synthetic division, (4x^4-4x^3+5x^2-4x+1)/(2x-1)